# Symplectic Lattices and GKP Codes - Simple Randomized Constructions from Cryptographic Lattices

### v1

@@ 12 Simulation of main results for ring constructions
In this section we show some simulation results in Section 10, i.e. the minimal distance for symplectic R-SIS and M-SIS lattices. The decoding performance for these lattices shows no difference from SIS lattices in Section 7, therefore we don’t put the results here.
@@
Figure 3 shows the simulation results of minimal distance for symplectic R-SIS and M-SIS lattices. For each combination of $(n,k,q)$, 1000 lattice samples are generated with $\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}$. Minimal distance of lattice is then calculated via BKZ reduction. In the figure, the red line represents the asymptotic bound $\sqrt{nk/\pi e}$ and the black dots represents the minimal distance for each lattice sample. The numbers at the bottom shows that for each combination $(n,k)$, how many samples have minimal distance $\lambda_{1}>\sqrt{nk/\pi e}$. Figure 3(a) corresponds to the case of Theorem 35, where $n=4$ is a power of 2, and $q=701$ satisfies $q=5$ mod $8$ and $q\geq\frac{\pi e(4\cdot 5)^{2}}{2}$. As shown in the figure, in all combinations of $(n,k)$, the generated M-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability almost to 1, which is even much larger then what the theorem tells; Figure 3(b) also has $n$ equals a power of 2, but a small $q=3$ is used in the simulation. Still the generated R-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability almost to 1. Thus from the simulation we make the conjecture that the generated M-SIS lattices have a high probability to achieve $\lambda_{1}\geq\sqrt{nk/\pi e}$ regardless of what $q$ is used for their generation. Figure 3(c) corresponds to the case of Theorem 36, where $n$ is a product of 2 primes. And for each $(n,k)$, certain $q$ equal to a primitive $\lambda$-root modulo $2n$ as well as $q\geq\frac{32\pi ek^{2}}{\ln(n/2)^{2}}$ is used in the simulation. The lattices still have higher probability to achieve $\lambda_{1}\geq\sqrt{nk/\pi e}$ than theorem suggested, but the probability is lower than the case where $n$ is a power of 2. This can relate to the factorization of polynomial $X^{n}+1$, which is mentioned in the proof of Theorem 36. Figure 3(d) corresponds to the case of Theorem 37, where $n$ is a product of 4 and a prime. Certain $q$ corresponds to the fist part of the theorem is used in the simulation. Again the results corresponds well with the theorem, and even suggest a higher probability of $\lambda_{1}\geq\sqrt{nk/\pi e}$.
@@ 13 Conclusion and Outlook
In this paper, we give the first efficient constructions of good families of GKP codes. Our constructions achieve the code distances non-constructively shown to be achievable by Preskill and Harrington [14] (based on results on symplectic lattices by Buser and Sarnak [19]) and conjectured by Conrad et al. to be achievable constructively by NTRU-based GKP codes [18]. Furthermore, the achieved distances are provably optimal or near optimal (up to small constants). Our GKP codes are constructed from well-known families of cryptographic lattices, SIS lattices. We also showed a trivial decoding algorithm for our GKP codes which does not use trapdoors. Experiments confirm the effectiveness of the constructions. For parameter choices and decoding algorithms the experiments even supersede the theoretical results.
@@ Acknowledgments
This work was funded by the Ministry of Culture and Science of the State of North Rhine-Westphalia and by the Federal Ministry of Research, Technology and Space (BMBFTR) (project PhoQuant). Additionally, this work was supported by the Equal Opportunity Program, Grant Line 2: Support for Female Junior Professors and Postdocs through Academic Staff Positions, 15th funding round of Paderborn University.

### v2

@@ 13 Simulation of main results for ring constructions
In this section we show some simulation results in Section 10, i.e. the minimal distance for symplectic R-SIS and M-SIS lattices. The decoding performance for these lattices shows no difference from SIS lattices in Section 7, so we omit those results here.
@@
Figure 3 shows the simulation results of minimal distance for symplectic R-SIS and M-SIS lattices. For each combination of $(n,k,q)$, 1000 lattice samples are generated with $\textbf{H}\leftarrow R_{q,k}^{\operatorname{Sym}}$. Minimal distance of lattice is then calculated via BKZ reduction. In the figure, the red line represents the asymptotic bound $\sqrt{nk/\pi e}$ and the black dots represent the minimal distance for each lattice sample. The numbers at the bottom show, for each combination $(n,k)$, how many samples have minimal distance $\lambda_{1}>\sqrt{nk/\pi e}$. Figure 3(a) corresponds to the case of Theorem 34, where $n=4$ is a power of 2, and $q=701$ satisfies $q=5\mod 8$ and $q\geq\frac{\pi e(4\cdot 5)^{2}}{2}$. As shown in the figure, in all combinations of $(n,k)$, the generated M-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability close to 1, which is even much larger than what the theorem tells; Figure 3(b) also has $n$ equal to a power of 2, but a small $q=3$ is used in the simulation. Still the generated R-SIS lattices have minimal distance larger than $\sqrt{nk/\pi e}$ with probability close to 1. Thus from the simulation we make the conjecture that the generated M-SIS lattices have a high probability of achieving $\lambda_{1}\geq\sqrt{nk/\pi e}$ regardless of what $q$ is used for their generation. Figures 3(c) and 3(d) illustrate more general instances covered by Theorem 35. In these examples, $n$ is not a power of 2, and the observed probabilities are lower than in the power-of-two case, which is consistent with the more delicate factorization behavior of $X^{n}+1$ modulo $q$. Again the results correspond well with the theorem, and even suggest a higher probability of $\lambda_{1}\geq\sqrt{nk/\pi e}$.
@@ 14 Conclusion and Outlook
In this paper, we give the first efficient constructions of good families of GKP codes. Our constructions achieve the code distances non-constructively shown to be achievable by Preskill and Harrington [18] (based on results on symplectic lattices by Buser and Sarnak [7]) and conjectured by Conrad et al. to be achievable constructively by NTRU-based GKP codes [11]. Furthermore, the achieved distances are provably optimal or near optimal (up to small constants). Our GKP codes are constructed from well-known families of cryptographic lattices, general SIS lattices, ring and module SIS lattices. Combining our results on GKP codes from R-SIS lattices with results from Stehlé and Steinfeld [32], we are able to prove Conjecture 2 from [11], namely, for certain parameter choices the NTRU-based construction of GKP codes yields GKP codes with distance $\sqrt{n/\pi e}$. We also showed a trivial decoding algorithm for our GKP codes which does not use trapdoors. Experiments confirm the effectiveness of the constructions. For parameter choices and decoding algorithms the experiments even supersede the theoretical results.
@@ Acknowledgments
This work was funded by the Ministry of Culture and Science of the State of North Rhine-Westphalia and by the German Federal Ministry of Research, Technology and Space (BMFTR) within the PhoQuant project (Grant No. 13N16103). Additionally, this work was supported by the Equal Opportunity Program, Grant Line 2: Support for Female Junior Professors and Postdocs through Academic Staff Positions, 15th funding round of Paderborn University.